OpenStudy (anonymous):

If 4^(2-x)=3 then 4^(2x-1)=?

OpenStudy (anonymous):

-x + 2 = log_4(3) 2x - 1 = log_4(?) x + 1 = log_4(3*?) 3*? = 4^(x+1) ? = (4^(x+1))/3

OpenStudy (anonymous):

should be a way to cancel x 1 minute

OpenStudy (anonymous):

-2x + 4 = log_4(9) 2x - 1 = log_4(y) 3 = log_4(9y) 9y = 64 y = 64/9

OpenStudy (anonymous):

how do you know where to add the log?

OpenStudy (anonymous):

which part?

OpenStudy (anonymous):

pls explain this part futher: -2x + 4 = log_4(9) 2x - 1 = log_4(y) 3 = log_4(9y)

OpenStudy (anonymous):

(4^(2-x))^2=3^2 4^(4-2x)=9 4^4=9*4^2x 4^3=9*4^(2x-1) 4^(2x-1)=64/9